OpenStudy (anonymous):

A = a+b+c / 3 Solve for b. A.) 3A - a - c = b B.) -c / 3 = b C.) A - c / 3 = b C.) -3 - c = b

OpenStudy (anonymous):

A

OpenStudy (agreene):

A

OpenStudy (anonymous):

A

OpenStudy (saifoo.khan):

A for Apple?

OpenStudy (anonymous):

A=a+b+c/3 3A=a+b+c 3A-a-c=b A.